OpenStudy (anonymous):

Simplify: (x - 8)^2. someone show meee step by step :)

OpenStudy (anonymous):

x^2 - 2 . x. 8 + 8^2 x^2 -16x +64

OpenStudy (anonymous):

(x - 8)^2 =x^2 -2(8)(x)+8^2 =x^2-16x+64

OpenStudy (anonymous):

Wowww !

OpenStudy (anonymous):

wowww for?

OpenStudy (anonymous):

Cuz im frustrated .

OpenStudy (anonymous):

x^2 - 2 . x. 8 + 8^2 x^2 -16x +64

OpenStudy (anonymous):

everyone is frustated about the field they are not interested in

OpenStudy (anonymous):

Thankss For understanding .

OpenStudy (anonymous):

x^2 - 16x + 64